# Self-Alignment: Improving Alignment of Cultural Values in LLMs via In-Context Learning

### v1

@@ Abstract
Introduction
The wide adoption of Large Language Models (LLMs) opened up pertinent questions as to how to correctly align LLM responses to reflect human intents and values (Cao et al. 2023; Shen et al. 2023; Liu et al. 2023; Xu et al. 2024; Wolf et al. 2023). While LLMs are already used by the public globally, much existing research shows that they are misaligned with respect to the cultural values that they encode (Arora, Kaffee, and Augenstein 2023; Choenni, Lauscher, and Shutova 2024) and tend to exhibit western-centric biases (Naous, Ryan, and Xu 2023). In practice though, to enable their adequate deployment in different languages, LLM output needs to be sensitive to the biases of the culturally diverse communities in which those languages are spoken (Hershcovich et al. 2022). Yet, popular alignment methods require large human preference datasets for fine-tuning and considerable computational resources (Rafailov et al. 2024; Ziegler et al. 2019), which makes it difficult and expensive to scale them to a multitude of languages and cultures. In this paper, we explore whether in-context learning (ICL), i.e., the LLMs’ ability to learn from a few demonstration examples at inference time (Wei et al. 2022; Brown et al. 2020; Dong et al. 2022), can be exploited to adjust the cultural values reflected in the LLM output, when provided with culturally-aligned demonstration examples. To the best of our knowledge, our work is the first to investigate this question.
@@ Introduction
Our demonstration examples are based on the questions from the World Values Survey (WVS) (Haerpfer et al. 2022), a social science effort documenting cultural values of participants in different countries. Specifically, we use the multilingual dataset of cultural value probes constructed by Arora, Kaffee, and Augenstein (2023) on the basis of the WVS. We use these probes both as prompts to probe the LLMs for their encoding of cultural values, and as our demonstration examples to evoke a cultural profile. Given that the probes are based on real survey questions, we can complete the demonstration examples with the real answer reported per country. See Figure 1 for an example.
@@
Related work
Misalignment of LLMs
While most popular LLMs have already undergone alignment in the form of reinforcement learning with human feedback (RLHF) (Ouyang et al. 2022) at the fine-tuning stage, various studies show that LLMs are still not adequately aligned to a wide range of human values. For instance, Santurkar et al. (2023) compared model opinions with human responses in public opinion polls among various demographic groups and found substantial positional misalignment. Durmus et al. (2023) expanded this study to a global scale using cross-national surveys and found a bias towards Western countries, as well as unwanted cultural stereotypes. He et al. (2024) instead studied affective alignment, and measured how the emotional and moral tone of LLMs represents those of different groups. Finally, Arora, Kaffee, and Augenstein (2023); Choenni, Lauscher, and Shutova (2024) and Cao et al. (2023) find that cultural values that LLMs encode in different languages do not align with human survey data, suggesting cultural misalignment.
Improving alignment at inference time
While there is a general consensus that LLMs should align to human values, such values vary considerably across countries, regions and even individuals. As such, it is underspecified what exactly we should aim to align the model to (Yao et al. 2023; Kirk et al. 2023). As a one-fits-all approach seems unlikely to lead to a satisfying outcome, and collecting a large alignment dataset for each scenario is prohibitively expensive, various approaches to improving alignment at inference time have been proposed. The benefit of such methods is that they would allow us to flexibly change alignment on an individual basis with minimal cost.
One such line of research focuses on sociodemographic prompting, a method to steer LLM responses towards answers that a persona, i.e., a human with a specific sociodemographic profile (e.g., age, gender, educational background, etc.), would give (Deshpande et al. 2023; Santurkar et al. 2023; Hwang, Majumder, and Tandon 2023; Cheng, Durmus, and Jurafsky 2023; He et al. 2024; Beck et al. 2024). While the methods used in these studies are also prompt-based, we exploit the ICL abilities of LLMs to trigger cultural profiles and instead improve alignment by inducing cultural knowledge into LLMs.
@@
We aim to improve the alignment of cultural values for different language/country combinations using demonstration examples constructed from the World Values Survey (WVS) data (Haerpfer et al. 2022). More specifically we use the cloze-style probing templates, created by Arora, Kaffee, and Augenstein (2023), based on the WVS data. The WVS collects data on cultural values in different countries in waves, and our questions come from Wave 7 which ran from 2017 to 2020 and targeted 57 countries 111https://www.worldvaluessurvey.org. Survey results are published per question, organised in 13 categories: (1) Social Values, Attitudes and Stereotypes, (2) Happiness and Well-being, (3) Social Capital, Trust and Organisational Membership, (4) Economic Values, (5) Corruption, (6) Migration, (7) Security, (8) Post-materialist Index, (9) Science and Technology, (10) Religious Values, (11) Ethical Values and Norms, (12) Political Interest and Political Participation, (13) Political Culture and Regimes. Categories (4) and (8) are excluded as their questions could not be converted into probes. We use 237 probes which always prompt the model to choose between two answers (e.g. important and unimportant or agree and disagree) for templates such as: ’Religion is _ to me.’ and ’I _ that when a woman works for pay, the children suffer.’.
@@
Arora, Kaffee, and Augenstein (2023) translated the English WVS probes into 13 languages: Romanian, Greek, Urdu, Farsi, Tagalog, Indonesian, German, Malay, Bengali, Serbian, Turkish, Vietnamese and Korean. These languages were selected using three criteria: (1) the languages can be mapped to one country covered by the WVS, (2) they are the official languages of the countries that they are mapped to, (3) the distribution of the language’s speakers can be primarily localized to a country or relatively small geographical region, and (4) all languages have at least 10K articles on Wikipedia such that the LLMs can have seen a sufficient amount of pretraining data. We use the multilingual probes to study how well our method performs in each language, when aligning to the cultural values of the country that this language is mapped to. For instance, we align LLMs in Romanian to the dominant responses for Romania from the WVS. See Appendix B for the full mapping.
Methods
Models
We test our method on 5 LLMs of varying sizes: Llama3-8B (Touvron et al. 2023), Mistral AI 7B (Jiang et al. 2023), CommandR 35B222https://docs.cohere.com/docs/command-r, Gemini-pro 1.5 50T333https://deepmind.google/technologies/gemini/ and BLOOMz 7B1 (Muennighoff et al. 2023). While most LLMs are English-centric, CommandR and BLOOMz were explicitly designed to be multilingual. For each LLM we use the instruction-tuned or chat fine-tuned version, see Appendix C for the full model details.
Prompt construction
We use the masked templates from Arora, Kaffee, and Augenstein (2023), but replace the [MASK] token with an underscore (see Figure 1 for an example of the prompt instructions). Moreover, we always complete each demonstration with the majority answer reported by the WVS results for the country that we are aligning to. Note that the WVS questions, in contrast to our probes, often ask for the degree to which the respondent agrees with the statement. As such, we aggregate the results from opposite ends of the scale to get a vote for each of our two classes (e.g. 1-5 is classified as disagree and 6-10 as agree). Finally, to ensure that the LLMs are not biased towards predicting one option over the other, we randomly pick which answer option to present first.
Demonstration selection strategies
@@ Methods
Besides selecting demonstration examples from the same WVS category, we use the chrF++ metric (Popović 2017) to determine the similarity between each test example and all possible demonstrations. ChrF++ calculates the character and word -gram overlap between two strings. As such, we use it as an inexpensive method to select demonstrations that have a greater lexical overlap with the test example (e.g. ’Friends are important’, ’Family is important’ etc.).
@@
Evaluation
Due to the stochastic nature of LLMs during the sampling process, generated LLM responses to the same prompt can vary. A common pitfall in the evaluation of LLMs is that in order to try to force deterministic outputs, researchers tend to set the temperature value to zero. This is not a satisfying solution for two reasons: (1) It does not allow for a realistic evaluation of LLMs as in practice users will likely not change the hyperparameters to enforce determinism hence resulting in a mismatch between the model under study and the one used in practice, and (2) contrary to popular belief, setting the temperature to zero does not always guarantee deterministic output (Ouyang et al. 2023). Thus, we instead embrace the stochastic nature of LLMs and use the default hyperparameters (see Appendix C for details), but evaluate a distribution of model responses.
@@
Self-alignment results in English
In Sections Differences in selection strategies and Robustness analysis, we first test which demonstration selection strategy is most effective using Llama3-8B in English when aligning to the values reported by the WVS survey for the United States. In Section Generalisation across models we then test how the best strategy performs across models.
Differences in selection strategies
@@
Generalisation across models
@@
Self-alignment in diverse languages
@@
Initial alignment to cultural values
@@
Before testing our method in different languages, we evaluate the initial alignment of the LLMs to the cultural values of the respective countries, in order to assess the degree of their misalignment. In Figure 7, we report the percentage of LLM responses that are misaligned across languages for each model/language combination. Note that we still apply strict criteria where examples that are not correctly answered across all 10 runs are classified as misaligned. We find that all models are relatively well aligned in English to US values compared to other language/country combinations. This is not surprising as LLMs are still predominantly trained on English data (Kew, Schottmann, and Sennrich 2023), and therefore tend to exhibit Western biases (Kotek, Dockum, and Sun 2023; Adilazuarda et al. 2024). Moreover, we find that overall, BLOOMz exhibits the worst alignment to human values and this result holds across all test languages. This could explain the lower effectiveness of self-alignment reported for BLOOMz in percentages in Section Generalisation across models, as there is a much larger number of misaligned examples in absolute numbers. Moreover, we find that BLOOMz also tends to be less consistent in its predictions, resulting in (almost) never being able to predict the right value 100 of the time. Moreover, across LLMs we find that test examples are especially poorly aligned in Romanian and Greek. For the distribution in error sizes across languages, see Appendix D. Overall, we find that these distributions are relatively similar across languages, except for CommandR.
Self-alignment results across languages
@@
We now use the multilingual probes described in Section Multilingual probes, and repeat the self-alignment experiments as before in different languages. The demonstration examples are selected using chrF++ from the dataset of the corresponding language. In Table 1, we find that self-alignment is effective across different languages. While the effectiveness on average drops slightly when using languages other than English, we also see many cases in which the method performs better on languages other than English. For instance, for Llama3-8B the method works best in Romanian (76), for CommandR in Vietnamese (72), for Gemini-pro in German (72), and for BLOOMz in Greek (69). Interestingly, we find that self-alignment is least biased to work well in English for BLOOMz (in 9/12 languages we are able to improve the alignment to a greater extent). This is interesting as it suggests that BLOOMz is not necessarily biased to adopt cultural values from the US more easily than for other countries. Importantly, BLOOMz is also the LLM for which the pretraining data was most balanced during training as it comprised of only 30 of English data. Moreover, in Section Initial alignment to cultural values, we found that for BLOOMz the highest number of examples were misaligned, meaning that in absolute numbers, self-alignment on BLOOMz does not underperform compared to the other LLMs.
| Demonstrations | Q&A | Explanation |
| I have no confidence in the press. I have no confidence in the banks. I have complete confidence in the courts. I have complete confidence in the police. I have no confidence in the parliament. | Q: I have _ confidence in the government. A: no | The pattern suggests that the speaker has no confidence in institutions that are typically associated with the executive branch (press, parliament), while having confidence in those that are part of the judicial and law enforcement branches (courts, police). |
| I agree that immigration increases employment. I disagree that immigration increases the crime rate. I disagree that immigration leads to social conflict. I agree that immigration strengthens cultural diversity. I agree that immigration fills important job vacancies | Q: I _ that immigration increases the risk of terrorism. A: disagree | Based on the previous statements, it seems that the individual has a generally positive view of immigration. Disagreeing that immi- gration increases the risk of terrorism is consistent with the previous statements, which showed a disagreement with negative views of immigration (crime rate, social conflict) and agreement with positive views (cultural diversity, filling job vacancies). |
| I agree that there is hell. I agree that there is God. I disagree that on the whole, men make better political leaders than women do. I disagree that on the whole, men make better business executives than women do. I agree that there is a heaven. | Q: On the whole, I _ that women are less corrupt than men. A: disagree | The reasoning is that the previous statements have a consistent pattern of: Agreeing with religious beliefs (hell, God, heaven). Disagreeing with gender-based generalizations (men vs. women in politics and business). So, it’s likely that the statement about women being less corrupt than men would also be disagreed with. |
@@
In previous sections, we quantitatively assessed the effectiveness of our method. We now perform a qualitative analysis to get a better idea of whether the LLM could be changing its predictions for the right reasons (Min et al. 2022). In particular, we repeat the experiments in English on misaligned examples, and this time, allow for more tokens to be generated such that the LLM can provide a local explanation to its generated answer post-hoc (Singh et al. 2024). In Table 2, we show three examples for which self-alignment led to improvements. In particular, we find that in most of such cases, the LLM explains its answer based on the patterns in the content of the demonstration examples. For instance, we find that the LLM distinguishes between confidence in different types of institutions (example 1), picks up on differences in attitudes towards positive and negative statements on immigration (example 2), and recognizes disagreement with gender-based generalizations and as such adjusts its responses to place men and women on the same footing (example 3).
@@
Despite these promising findings, we also in some (rare) cases find that the model adjusts its responses based on spurious correlations (e.g., ‘Based on the previous examples, it seems that the pattern is to agree with the statement if previous statements were also agreed with, and disagree if previous statements were disagreed with.’). Future work should study (1) how faithful these generated explanations are (Chen et al. 2023; Ye and Durrett 2022), and (2) how we can avoid the tendency of LLMs to exploit superficial patterns in the demonstration examples. In particular, this could reduce the cases for which we observe that alignment deteriorates instead of improves due to self-alignment.
Conclusion

### v2

@@ Abstract
1 Introduction
The wide adoption of Large Language Models (LLMs) opened up pertinent questions as to how to correctly align LLM responses to reflect human intents and values (Cao et al., 2023; Shen et al., 2023; Liu et al., 2023; Xu et al., 2024; Wolf et al., 2023). While LLMs are already used by the public globally, much existing research shows that they are misaligned with respect to the cultural values that they encode (Arora et al., 2023; Choenni et al., 2024) and tend to exhibit western-centric biases (Naous et al., 2023). In practice though, to enable their adequate deployment in different languages, LLM output needs to be sensitive to the biases of the culturally diverse communities in which those languages are spoken (Hershcovich et al., 2022). Yet, popular alignment methods require large human preference datasets for fine-tuning and considerable computational resources (Rafailov et al., 2024; Ziegler et al., 2019), which makes it difficult and expensive to scale them to a multitude of languages and cultures. In this paper, we explore whether in-context learning (ICL), i.e., the LLMs’ ability to learn from a few demonstration examples at inference time (Wei et al., 2022; Brown et al., 2020; Dong et al., 2022), can be exploited to adjust the cultural values reflected in the LLM output, when provided with culturally-aligned demonstration examples. To the best of our knowledge, our work is the first to investigate this question.
@@ 1 Introduction
Our demonstration examples are based on the questions from the World Values Survey (WVS) (Haerpfer et al., 2022), a social science effort documenting cultural values of participants in different countries. Specifically, we use the multilingual dataset of cultural value probes constructed by Arora et al. (2023) on the basis of the WVS. We use these probes both as prompts to probe the LLMs for their encoding of cultural values, and as our demonstration examples to evoke a cultural profile. Given that the probes are based on real survey questions, we can complete the demonstration examples with the real answer reported per country. See Figure 1 for an example.
@@
2 Related work
2.1 Misalignment of LLMs
While most popular LLMs have already undergone alignment in the form of reinforcement learning with human feedback (RLHF) (Ouyang et al., 2022) at the fine-tuning stage, various studies show that LLMs are still not adequately aligned to a wide range of human values. For instance, Santurkar et al. (2023) compared model opinions with human responses in public opinion polls among various demographic groups and found substantial positional misalignment. Durmus et al. (2023) expanded this study to a global scale using cross-national surveys and found a bias towards Western countries, as well as unwanted cultural stereotypes. He et al. (2024) instead studied affective alignment, and measured how the emotional and moral tone of LLMs represents those of different groups. Finally, Arora et al. (2023); Choenni et al. (2024) and Cao et al. (2023) find that cultural values that LLMs encode in different languages do not align with human survey data, suggesting cultural misalignment.
2.2 Improving alignment at inference time
While there is a general consensus that LLMs should align to human values, such values vary considerably across countries, regions and even individuals. As such, it is underspecified what exactly we should aim to align the model to (Yao et al., 2023; Kirk et al., 2023). As a one-fits-all approach seems unlikely to lead to a satisfying outcome, and collecting a large alignment dataset for each scenario is prohibitively expensive, various approaches to improving alignment at inference time have been proposed. The benefit of such methods is that they would allow us to flexibly change alignment on an individual basis with minimal cost.
One such line of research focuses on sociodemographic prompting, a method to steer LLM responses towards answers that a persona, i.e., a human with a specific sociodemographic profile (e.g., age, gender, educational background, etc.), would give (Deshpande et al., 2023; Santurkar et al., 2023; Hwang et al., 2023; Cheng et al., 2023; He et al., 2024; Beck et al., 2024). While the methods used in these studies are also prompt-based, we exploit the ICL abilities of LLMs to trigger cultural profiles and instead improve alignment by inducing cultural knowledge into LLMs.
@@ 3 Dataset
We aim to improve the alignment of cultural values for different language/country combinations using demonstration examples constructed from the World Values Survey (WVS) data (Haerpfer et al., 2022). More specifically we use the cloze-style probing templates, created by Arora et al. (2023), based on the WVS data. The WVS collects data on cultural values in different countries in waves, and our questions come from Wave 7 which ran from 2017 to 2020 and targeted 57 countries 111https://www.worldvaluessurvey.org. Survey results are published per question, organised in 13 categories: (1) Social Values, Attitudes and Stereotypes, (2) Happiness and Well-being, (3) Social Capital, Trust and Organisational Membership, (4) Economic Values, (5) Corruption, (6) Migration, (7) Security, (8) Post-materialist Index, (9) Science and Technology, (10) Religious Values, (11) Ethical Values and Norms, (12) Political Interest and Political Participation, (13) Political Culture and Regimes. Categories (4) and (8) are excluded as their questions could not be converted into probes. We use 237 probes which always prompt the model to choose between two answers (e.g. important and unimportant or agree and disagree) for templates such as: ’Religion is _ to me.’ and ’I _ that when a woman works for pay, the children suffer.’.
@@
Arora et al. (2023) translated the English WVS probes into 13 languages: Romanian, Greek, Urdu, Farsi, Tagalog, Indonesian, German, Malay, Bengali, Serbian, Turkish, Vietnamese and Korean. These languages were selected using three criteria: (1) the languages can be mapped to one country covered by the WVS, (2) they are the official languages of the countries that they are mapped to, (3) the distribution of the language’s speakers can be primarily localized to a country or relatively small geographical region, and (4) all languages have at least 10K articles on Wikipedia such that the LLMs can have seen a sufficient amount of pretraining data. We use the multilingual probes to study how well our method performs in each language, when aligning to the cultural values of the country that this language is mapped to. For instance, we align LLMs in Romanian to the dominant responses for Romania from the WVS. See Appendix B for the full mapping.
4 Methods
4.1 Models
We test our method on 5 LLMs of varying sizes: Llama3-8B (Touvron et al., 2023), Mistral AI 7B (Jiang et al., 2023), CommandR 35B222https://docs.cohere.com/docs/command-r, Gemini-pro 1.5 50T333https://deepmind.google/technologies/gemini/ and BLOOMz 7B1 (Muennighoff et al., 2023). While most LLMs are English-centric, CommandR and BLOOMz were explicitly designed to be multilingual. For each LLM we use the instruction-tuned or chat fine-tuned version, see Appendix C for the full model details.
4.2 Prompt construction
We use the masked templates from Arora et al. (2023), but replace the [MASK] token with an underscore (see Figure 1 for an example of the prompt instructions). Moreover, we always complete each demonstration with the majority answer reported by the WVS results for the country that we are aligning to. Note that the WVS questions, in contrast to our probes, often ask for the degree to which the respondent agrees with the statement. As such, we aggregate the results from opposite ends of the scale to get a vote for each of our two classes (e.g. 1-5 is classified as disagree and 6-10 as agree). Finally, to ensure that the LLMs are not biased towards predicting one option over the other, we randomly pick which answer option to present first.
4.3 Demonstration selection strategies
@@
Besides selecting demonstration examples from the same WVS category, we use the chrF++ metric Popović (2017) to determine the similarity between each test example and all possible demonstrations. ChrF++ calculates the character and word -gram overlap between two strings. As such, we use it as an inexpensive method to select demonstrations that have a greater lexical overlap with the test example (e.g. ’Friends are important’, ’Family is important’ etc.).
@@
4.4 Evaluation
Due to the stochastic nature of LLMs during the sampling process, generated LLM responses to the same prompt can vary. A common pitfall in the evaluation of LLMs is that in order to try to force deterministic outputs, researchers tend to set the temperature value to zero. This is not a satisfying solution for two reasons: (1) It does not allow for a realistic evaluation of LLMs as in practice users will likely not change the hyperparameters to enforce determinism hence resulting in a mismatch between the model under study and the one used in practice, and (2) contrary to popular belief, setting the temperature to zero does not always guarantee deterministic output (Ouyang et al., 2023). Thus, we instead embrace the stochastic nature of LLMs and use the default hyperparameters (see Appendix C for details), but evaluate a distribution of model responses.
@@ 4.5 Alignment procedure
5 Self-alignment results in English
In Sections 5.1 and 5.2, we first test which demonstration selection strategy is most effective using Llama3-8B in English when aligning to the values reported by the WVS survey for the United States. In Section 5.3 we then test how the best strategy performs across models.
5.1 Differences in selection strategies
@@ 5.2 Robustness analysis
5.3 Generalisation across models
@@
6 Self-alignment in diverse languages
@@
6.1 Initial alignment to cultural values
@@
Before testing our method in different languages, we evaluate the initial alignment of the LLMs to the cultural values of the respective countries, in order to assess the degree of their misalignment. In Figure 7, we report the percentage of LLM responses that are misaligned across languages for each model/language combination. Note that we still apply strict criteria where examples that are not correctly answered across all 10 runs are classified as misaligned. We find that all models are relatively well aligned in English to US values compared to other language/country combinations. This is not surprising as LLMs are still predominantly trained on English data (Kew et al., 2023), and therefore tend to exhibit Western biases (Kotek et al., 2023; Adilazuarda et al., 2024). Moreover, we find that overall, BLOOMz exhibits the worst alignment to human values and this result holds across all test languages. This could explain the lower effectiveness of self-alignment reported for BLOOMz in percentages in Section 5.3, as there is a much larger number of misaligned examples in absolute numbers. Moreover, we find that BLOOMz also tends to be less consistent in its predictions, resulting in (almost) never being able to predict the right value 100 of the time. Moreover, across LLMs we find that test examples are especially poorly aligned in Romanian and Greek. For the distribution in error sizes across languages, see Appendix D. Overall, we find that these distributions are relatively similar across languages, except for CommandR.
6.2 Self-alignment results across languages
@@
We now use the multilingual probes described in Section 3, and repeat the self-alignment experiments as before in different languages. The demonstration examples are selected using chrF++ from the dataset of the corresponding language. In Table 1, we find that self-alignment is effective across different languages. While the effectiveness on average drops slightly when using languages other than English, we also see many cases in which the method performs better on languages other than English. For instance, for Llama3-8B the method works best in Romanian (76), for CommandR in Vietnamese (72), for Gemini-pro in German (72), and for BLOOMz in Greek (69). Interestingly, we find that self-alignment is least biased to work well in English for BLOOMz (in 9/12 languages we are able to improve the alignment to a greater extent). This is interesting as it suggests that BLOOMz is not necessarily biased to adopt cultural values from the US more easily than for other countries. Importantly, BLOOMz is also the LLM for which the pretraining data was most balanced during training as it comprised of only 30 of English data. Moreover, in Section 6.1, we found that for BLOOMz the highest number of examples were misaligned, meaning that in absolute numbers, self-alignment on BLOOMz does not underperform compared to the other LLMs.
| Demonstrations | Q&A | Explanation | |||||||||||||||||
@@ 7 Qualitative analysis
In previous sections, we quantitatively assessed the effectiveness of our method. We now perform a qualitative analysis to get a better idea of whether the LLM could be changing its predictions for the right reasons (Min et al., 2022). In particular, we repeat the experiments in English on misaligned examples, and this time, allow for more tokens to be generated such that the LLM can provide a local explanation to its generated answer post-hoc (Singh et al., 2024). In Table 2, we show three examples for which self-alignment led to improvements. In particular, we find that in most of such cases, the LLM explains its answer based on the patterns in the content of the demonstration examples. For instance, we find that the LLM distinguishes between confidence in different types of institutions (example 1), picks up on differences in attitudes towards positive and negative statements on immigration (example 2), and recognizes disagreement with gender-based generalizations and as such adjusts its responses to place men and women on the same footing (example 3).
@@
Despite these promising findings, we also in some (rare) cases find that the model adjusts its responses based on spurious correlations (e.g., ‘Based on the previous examples, it seems that the pattern is to agree with the statement if previous statements were also agreed with, and disagree if previous statements were disagreed with.’). Future work should study (1) how faithful these generated explanations are (Chen et al., 2023; Ye and Durrett, 2022), and (2) how we can avoid the tendency of LLMs to exploit superficial patterns in the demonstration examples. In particular, this could reduce the cases for which we observe that alignment deteriorates instead of improves due to self-alignment.
8 Conclusion